OpenStudy (anonymous):

Please help? I don't know how to solve for concentration: If it requires 25.0 milliliters of 1.0 molar HCl to neutralize 45.0 milliliters of NaOH, what is the concentration of the NaOH solution? Balanced equation: NaOH + HCl yields NaCl + H2O 0.56 M NaOH 0.82 M NaOH 1.1 M NaOH 1.8 M NaOH

OpenStudy (anonymous):

Would it be something like: n = c x v = 1 x 0.025 = 0.025 moles = n NaOH = n/v = 0.025/0.045 = 0.56M So is it A? 0.56 M NaOH?

OpenStudy (jfraser):

That's the way to do it

OpenStudy (sweetburger):

Did you use M1V1=M2V2?